Circular No. 5624

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET GEHRELS 3 (1992v) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the Spacewatch telescope at Kitt Peak. The object is of stellar appearance and the position is in close agreement with the prediction on MPC 16381 (ephemeris on MPC 20171).

1992 UT  α_{2000}  δ_{2000}  m_{1}
Sept.26.44826  5^{h}25^{m}12.^{s}80  +23^{o}08′19′′.1  22.0
26.45819  5 25 12.99  +23 08 19.8
28.45122  5 26 02.94  +23 08 35.5
28.46007  5 26 03.06  +23 08 35.3
28.47270  5 26 03.43  +23 08 35.7  21.6

SATURN VI (TITAN) M. T. Lemmon and E. Karkoschka, Lunar and Planetary Laboratory,

University of Arizona, report: “We observed Titan on July 18 and 19 at the Steward Observatory 2.3-m telescope and on Sept. 13, 14, 15, and 17 at the Multiple Mirror Telescope, using the Ge Spectrograph (range 0.85-1.35 μm, resolution 0.01 μm; calibration done relative to the nearby star SAO 164338). While Titan’s spectral shape remained constant throughout our observing periods in regions of medium and strong methane absorption, we see significant variations in continuum wavelengths. In the continuum regions of 1.08 and 1.28 μm, Titan is 14 and 22 percent brighter, respectively, on Sept. 13 and 14 (when it was near eastern elongation) with respect to July 18 and 19 (near western elongation). A small observed brightness increase at 0.94 μm is less certain, due to interference from telluric absorption. The continuum albedos for Sept. 15 are below those for Sept. 13 and 14; by Sept. 17, when only the 1.28-μm region was observed, the continuum albedo was consistent with the July observations. Our atmospheric models of Titan indicate that the observed spectral variations derive from clouds below 10 km altitude, or, more likely, from the surface. If we indeed see the surface, we can constrain surface albedo variations and Titan’s rotational period. At 1.08 and 1.28 μm, Titan has an average surface albedo near 0.2, and the minimum to maximum variations are at least 0.1. Our data are consistent with a synchronous rotational period of 16 days, but can also be fit with certain rotational periods from 5 to 18 days.”

1992 September 28 Daniel W. E. Green